OpenStudy (anonymous):

find the following ^14sqrt(-8)^14

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[14]{(-8)^{14}} = |-8| = 8\] So \[\Large \sqrt[14]{(-8)^{14}} = 8\]

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

I used the idea that \[\Large \sqrt[n]{x^n}=|x|\] where n is an even integer

OpenStudy (anonymous):

what do you do witht the 14 i mean ^ 14 nothing

jimthompson5910 (jim_thompson5910):

The 14's effectively cancel out

OpenStudy (anonymous):

oh alright !thanks